# Transmission lines; Defect detection; UAV imagery; Power industry; Transformer-based detectors

### v2

@@ 1 INTRODUCTION
With the rapid advancement of smart grid development, transmission line defect detection has increasingly adopted deep learning-based object detection methodsLi et al. (2023). Alongside this progress, numerous general-purpose detectors—such as YOLO(Redmon et al., 2016), DINOZhang et al. (2022), and RT-DETR(Zhao et al., 2024)—have been developed. While these models achieve remarkable success in generic scenarios, they face notable limitations in transmission line defect detection under complex conditions involving occlusion, background clutter, and variable lightingLiu et al. (2023). Common issues include missed detections, false positives, and false negatives, which substantially undermine the practical utility of UAV-acquired imagery in real-world transmission line maintenanceDeng et al. (2024). These shortcomings are consistently reported in previous studies and further confirmed by our preliminary experiments, highlighting the critical need for specialized detection frameworks capable of accurately identifying small-scale defects in UAV operational scenarios.
@@
    We design a novel Edge-Enhanced Convolution (EEConv) module that integrates central, horizontal, and vertical difference convolutions to embed gradient priors into feature extraction. This enhances sensitivity to fine-grained details and object boundaries. Through a re-parameterization strategy, the multi-branch structure is fused into a single $3\times 3$ kernel at inference, ensuring no extra computational overhead while substantially improving the representation of small-scale defects in UAV inspection imagery.
@@ 3.1 Overview
In this study, RE-DETR-r18 is employed as the baseline detection framework owing to its superior accuracy and robustness in complex scenarios. Traditional DETR models often suffer from slow convergence and substantial computational overhead, which hinder their deployment in UAV-based inspection tasks. Addressing these inherent limitations, RE-DETR incorporates an enhanced encoder–decoder architecture together with refined query interaction mechanisms, while for the backbone network, RT-DETR-r18 adopts a CNN-based design, namely the popular ResNet 18. From the perspective of real-time performance, employing a CNN architecture for feature extraction further enhances the real-time capability of DETR-like models, thereby improving their practical applicability. As a result, RE-DETR not only accelerates convergence and improves detection precision, but also preserves the global context modeling capability of Transformer-based frameworks(Vaswani et al., 2017).In addition, its lightweight optimization strategies—specifically, the design choice to apply the encoder only to high-level semantic features (S5) with relatively small spatial dimensions—effectively reduce resource consumption. This balances performance and speed by significantly decreasing computational overhead and accelerating inference without compromising accuracy, thereby making RE-DETR particularly suitable for edge deployment in UAV applications characterized by stringent computational and energy constraints(Tian et al., 2025).
@@
By contrast, YOLO-based detectors, though widely favored for their real-time efficiency and ease of deployment(Vijayakumar and Vairavasundaram, 2024), tend to degrade in performance when handling small, occluded, or densely distributed objects, owing to their anchor-dependent design and limited global receptive field. This aligns with the findings of Ismail et al.(Ismail et al., 2024), who reported that while YOLO excels in real-time scenarios, it struggles with small-object detection and occlusion, whereas DETR achieves higher accuracy in complex environments at the cost of greater computational demand. Our preliminary experiments further compared RT-DETR with representative YOLO series models, including YOLOv3(Redmon and Farhadi, 2018), YOLOv7(Wang et al., 2022), and YOLOv11(Khanam and Hussain, 2024), and showed that RT-DETR consistently outperformed most of them in overall performance. In cases where certain YOLO variants (e.g., YOLO 11m) surpassed RT-DETR, their gains were typically accompanied by substantially larger parameter sizes, thereby underscoring RT-DETR’s advantageous balance between model complexity and detection accuracy. Taken together, the consistent advantages of RT-DETR over both traditional DETR models and representative YOLO series substantiate its adoption as the baseline framework in this study. Building upon this strong foundation, we propose TinyDef-DETR as a dedicated solution for small object defect detection in UAV-based transmission line inspection. The overall architecture is illustrated in Fig.1.
@@
First, the backbone adopts the Detail Enhanced ResNet (DE-ResNet) structure to strengthen low-level detail preservation. Second, a compact Space-to-Depth Convolution (S2DConv) module is inserted before each downsampling layer to reorganize spatial features in a manner that preserves critical spatial information while reducing resolution, thereby improving the recall rate of extremely small objects. Third, a lightweight Cross-Stage Partial Attention Block (CSPAB) is embedded into the backbone network to enhance directional sensitivity and fine-grained structural modeling, effectively capturing anisotropic features of typical defects such as missing bolts, loose clamps, or sagging wires. Detailed descriptions of the proposed modules can be found in Sections 3.2 to  3.4.
@@ 3.2 Edge Enhanced ResNet
ResNet(He et al., 2016) represents a milestone in deep learning for computer vision, introducing residual learning that enables the effective training of very deep networks. Owing to its simple yet modular design, it has become a widely adopted backbone across a broad spectrum of vision tasks. Nevertheless, when applied to UAV-based visual inspection for transmission line defect detection, native ResNet architectures, as verified in preliminary experiments, exhibit limited capability in capturing sufficiently fine-grained features, thereby constraining their effectiveness in identifying small and subtle defect targets, particularly in challenging transmission line inspection environments with interference from vegetation, man-made structures, and adjacent electrical installations. Inspection of the heatmaps exported from ResNet further reveals that they lack clear edges and fine contours, which exacerbates the difficulty of accurately localizing subtle defects. Unlike conventional convolution, where the high performance of CNN-based edge feature extraction typically relies on large pretrained backbones—resulting in high memory consumption and energy cost, which contradicts the real-time requirements of UAV-based transmission line defect detection tasks—difference convolution introduces differential operations to enhance sensitivity to gradient variations and fine-grained structural details, thereby improving the extraction of local details and object boundariesWang et al. (2025). To address the aforementioned problem, we propose an EEConv (edge enhancement convolution). Building on this idea, EEConv explicitly incorporates multiple difference convolution operators—such as central difference convolution (CDC), horizontal difference convolution (HDC), and vertical difference convolution (VDC)—to embed gradient priors into the feature extraction process. The overall structure of EEConv is illustrated in Fig. 2a.To ensure deployment efficiency, EEConv adopts a re-parameterization strategy that fuses its multi-branch convolutions into a single equivalent 3×3 kernel during inference. Once fused, the auxiliary branches are discarded, leaving only a lightweight standard convolution. This allows the module to preserve the expressive power gained from multi-branch training while maintaining the same computational cost as a vanilla convolution.During deployment, the structure of EEConv becomes as shown in Fig. 2b.Building upon EEConv, we further design an Edge Enhancement Block (EEBlock), which stacks multiple EEConvs together and integrates normalization and non-linear activation to form a complete feature extraction unit, as depicted in Fig. 2c.
@@
Formally, let the four convolutional branches be denoted as $(W_{k},b_{k}),\;k=1,\dots,4$, corresponding to CDC, HDC, VDC and the conventional convolution, respectively. During training, their weights and biases are linearly combined:
$W_{\Sigma}$ $=\sum_{k=1}^{4}W_{k},$ (1)
$b_{\Sigma}$ $=\sum_{k=1}^{4}b_{k}.$ (2)
@@
To integrate this into our framework, we replace the second $3\times 3$ convolution in the residual branch of standard ResNet basic blocks with EECConv, thereby preserving the backbone’s modular design while substantially improving sensitivity to fine-grained structures.
This design choice is motivated by the observation that the second $3\times 3$ convolution in each residual block primarily functions as a feature refiner, responsible for aggregating local structural details after the first convolution has extracted preliminary features. In this way, we obtain the EEBlock, whose structure is illustrated in Fig.LABEL:EEBlock
By substituting it with EEConv, we enable the block to embed directional gradient priors precisely at the stage where fine-grained information is consolidated, thus maximizing the impact of EEConv on enhancing edge and detail representation. Meanwhile, the first $3\times 3$ convolution and the shortcut connection remain unchanged, ensuring compatibility with the original ResNet architecture and avoiding disruption of its residual learning mechanism.
By embedding EEConv into the backbone, TinyDef-DETR achieves improved sensitivity to subtle structural details without introducing additional inference overhead. This enhancement directly addresses the challenge of detecting small-scale defects in transmission line imagery, such as polluted or broken insulators, damaged poles, missing or loose tie wires, insulator flashovers, and even bird nests. Moreover, the integration of directional gradient priors strengthens boundary-aware features, leading to higher localization accuracy and improved robustness against background interference. Importantly, the efficiency of the re-parameterized form ensures that TinyDef-DETR remains lightweight, scalable, and practical for large-scale, real-time UAV-based inspection scenarios.
@@ 3.4 Cross-Stage Dual-Domain Multi-Scale Attention Module
In the task of transmission line defect detection, most targets are small objects such as cracks, rust spots, or insulator damages. These defects usually exhibit weak visual saliency and subtle structural variations, which makes them difficult to distinguish from complex backgrounds including vegetation, sky, or towers. Standard convolutional networks often struggle in such cases because local convolutions have limited receptive fields and cannot effectively capture global dependencies, while large kernels may introduce high computational costs and oversmoothing. Similarly, conventional attention mechanisms that focus only on spatial or channel domains fail to provide sufficient enhancement for small, noisy targets. Therefore, a more effective solution is required that can simultaneously preserve fine-grained local details, capture global semantic context, and adaptively highlight defect-related features across both frequency and spatial domains.
To meet these requirements, we propose the Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM). Given input features $X\in\mathbb{R}^{B\times C\times H\times W}$, a $1\times 1$ convolution is first applied to adjust the channel dimension. The resulting features are then processed in parallel by a large branch, a local branch, and a dual-domain attention branch (FCSA), and the outputs are aggregated through residual summation and further modulated by a $1\times 1$ convolution.
In the large branch, we employ three types of depth-wise convolutions with kernel sizes $(k,k),(1,k),(k,1)$, where $k=31$, to capture multi-directional large-scale receptive fields with minimal computational overhead. We employ the square kernel to model holistic global context, while we employ the strip-shaped kernels to extract elongated dependencies along horizontal and vertical directions. This configuration allows the network to better perceive line-shaped structures, which are typical in transmission line defects.
In the local branch, we employ a simple $1\times 1$ depth-wise convolution to preserve pixel-level details that are critical for detecting small-scale defects. This operation ensures that fine-grained cues, such as cracks or corrosion spots, are maintained and not smoothed out by large-kernel operations.
In the dual-domain attention branch, we introduce the Frequency-Channel-Spatial Attention (FCSA) mechanism, which enhances feature representation by integrating frequency-domain filtering and spatial-channel reweighting into a unified formulation. Given the input feature map $X\in\mathbb{R}^{B\times C\times H\times W}$, we first apply the two-dimensional Fourier transform to obtain the frequency representation:
@@
where $\mathcal{F}$ denotes the two-dimensional discrete Fourier transform (DFT). To emphasize informative frequency components, we generate a set of channel-wise weights $A_{c}\in\mathbb{R}^{B\times C\times 1}$ through global average pooling followed by a $1\times 1$ convolution. The frequency-domain enhancement is then expressed as:
$\tilde{X}_{f}=A_{c}\cdot X_{f}.$
We subsequently reconstruct the enhanced spatial representation by applying the inverse Fourier transform:
$X_{fca}=\mathcal{F}^{-1}(\tilde{X}_{f}),$
where $\mathcal{F}^{-1}$ denotes the two-dimensional inverse Fourier transform (IDFT). On top of this frequency enhancement, we compute a spatial-channel attention map $A_{sc}\in\mathbb{R}^{B\times C\times 1}$, which is obtained from $X_{fca}$ using global average pooling and a $1\times 1$ convolution. This map adaptively reweights the features in the spatial domain:
$X_{fca}=A_{sc}\cdot X_{fca}.$
Finally, the refined representation $X_{fca}$ is further calibrated by a lightweight Feature Guidance Module (FGM) before being integrated with the other branches. In this way, FCSA selectively enhances defect-related features by jointly exploiting frequency filtering, channel emphasis, and spatial reweighting, thereby improving the discriminability of small targets in complex backgrounds.
The outputs of all branches are integrated to form the final response of the module:
$Y=X+DW_{k,k}(X_{n})+DW_{k,1}(X_{n})+DW_{k,k}(X_{n})+DW_{1,1}(X_{n})+X_{fca}.$
To improve efficiency and gradient flow, we extend the design into a cross-stage structure, referred to as CSDMAM. Following the Cross Stage Partial (CSP) strategy, we divide the input into two parts: one part is processed by the proposed multi-branch attention module, while the other part is preserved as an identity mapping. We then concatenate the two parts and employ a $1\times 1$ convolution to fuse them, yielding:
$Y_{cs}=\text{Conv}([\text{CSDMAM}(X_{\text{branch}}),X_{\text{identity}}]).$
By combining multi-scale receptive fields, local detail preservation, and the unified FCSA dual-domain attention mechanism, we employ CSDMAM to achieve a balance between global semantic understanding and fine-grained localization. With the cross-stage integration strategy, we further improve computational efficiency and training stability. As a result, CSDMAM is particularly effective for transmission line defect detection, where most targets are small, visually ambiguous, and require both robust context modeling and precise localization.
Figure 4: The architecture of Cross-Stage Dual-Domain Multi-Scale Attention Module

### v3

@@ 1 INTRODUCTION
With the rapid advancement of smart grid development, transmission line defect detection has increasingly adopted deep learning-based object detection methodsLi et al. (2023). Alongside this progress, numerous general-purpose detectors—such as YOLO(Redmon et al., 2016), DINOVaswani et al. (2017), and RT-DETR(Zhao et al., 2024)—have been developed. While these models achieve remarkable success in generic scenarios, most of them on the COCO datasetLin et al. (2014), they face notable limitations in transmission line defect detection under complex conditions involving occlusion, background clutter, and variable lightingLiu et al. (2023). Common issues include missed detections, false positives, and false negatives, which substantially undermine the practical utility of UAV-acquired imagery in real-world transmission line maintenanceDeng et al. (2024). These shortcomings are consistently reported in previous studies and further confirmed by our preliminary experiments, highlighting the critical need for specialized detection frameworks capable of accurately identifying small-scale defects in UAV operational scenarios.
@@
    We construct an Edge Enhanced ResNet (EE-ResNet) by embedding an Edge-Enhanced Convolution (EEConv) module into the residual backbone. The EEConv integrates central, horizontal, and vertical difference convolutions to encode gradient priors, thereby improving sensitivity to fine-grained details and object boundaries. Through a re-parameterization strategy, the multi-branch structure of EEConv is fused into a single $3\times 3$ kernel at inference, incurring no additional computational overhead. By augmenting the ResNet backbone with such edge-aware feature extraction, EE-ResNet substantially strengthens the representation capability for small-scale defects in UAV inspection imagery.
@@ 3.1 Overview
In this study, we employ RE-DETR-r18Zhao et al. (2024) as the baseline detection framework owing to its superior accuracy and robustness in complex scenarios. Traditional DETR models often suffer from slow convergence and substantial computational overhead, which hinder their deployment in UAV-based inspection tasks.
@@
Addressing these inherent limitations, RE-DETR incorporates an enhanced encoder–decoder architecture together with refined query interaction mechanisms, while for the backbone network, RT-DETR-r18 adopts a CNN-based design, namely the popular ResNet-18He et al. (2016). From the perspective of real-time performance, employing a CNN architecture for feature extraction further enhances the real-time capability of DETR-like models, thereby improving their practical applicability. As a result, RE-DETR not only accelerates convergence and improves detection precision, but also preserves the global context modeling capability of Transformer-based frameworks(Vaswani et al., 2017).In addition, its lightweight optimization strategies—specifically, the design choice to apply the encoder only to high-level semantic features (S5) with relatively small spatial dimensions—effectively reduce resource consumption. This balances performance and speed by significantly decreasing computational overhead and accelerating inference without compromising accuracy, thereby making RE-DETR particularly suitable for post-flight UAV imagery analysis, where large volumes of data must be processed efficiently under computational and energy constraints(Tian et al., 2025).
@@
By contrast, YOLO-based detectors, though widely favored for their real-time efficiency and ease of deployment(Vijayakumar and Vairavasundaram, 2024), tend to degrade in performance when handling small, occluded, or densely distributed objects, owing to their anchor-dependent design and limited global receptive field. This aligns with the findings of Ismail et al.(Ismail et al., 2024), who reported that while YOLO excels in real-time scenarios, it struggles with small-object detection and occlusion, whereas DETR achieves higher accuracy in complex environments at the cost of greater computational demand. Our preliminary experiments further compared RT-DETR with representative YOLO series models, including YOLOv3(Redmon and Farhadi, 2018), YOLOv7(Wang et al., 2023), and YOLO 11(Khanam and Hussain, 2024), and showed that RT-DETR consistently outperformed most of them in overall performance. In cases where certain YOLO variants (e.g., YOLO 11m) surpassed RT-DETR, their gains were typically accompanied by substantially larger parameter sizes, thereby underscoring RT-DETR’s advantageous balance between model complexity and detection accuracy.
Taken together, the consistent advantages of RT-DETR over both traditional DETR models and representative YOLO series substantiate its adoption as the baseline framework in this study.
Building upon this strong foundation, we propose TinyDef-DETR as a dedicated solution for small object defect detection in UAV-based transmission line inspection. The overall architecture is illustrated in Fig.1.
@@
First, the backbone adopts the Edge-Enhanced ResNet (EE-ResNet) structure to strengthen low-level detail preservation. Second, a compact Space-to-Depth Convolution (S2DConv) module is inserted before each downsampling layer to reorganize spatial features in a manner that preserves critical spatial information while reducing resolution, thereby improving the recall rate of extremely small objects. Third, a lightweight Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM) is embedded into the backbone network to enhance directional sensitivity and fine-grained structural modeling, effectively capturing anisotropic features of typical defects such as missing bolts, loose clamps, or sagging wires. Finally, a Focaler-Wise-SIoU loss function is employed to jointly optimize classification and localization. This loss function adaptively balances easy and hard samples while providing more accurate bounding box regression through shape-aware optimization, which further improves detection robustness for small and ambiguous targets. Detailed descriptions of the proposed modules can be found in Sections 3.2 to  3.5.
@@ 3.2 Edge Enhanced ResNet
ResNet(He et al., 2016) represents a milestone in deep learning for computer vision, introducing residual learning that enables the effective training of very deep networks. Owing to its simple yet modular design, it has become a widely adopted backbone across a broad spectrum of vision tasks. Nevertheless, when applied to UAV-based visual inspection for transmission line defect detection, native ResNet architectures, as verified in preliminary experiments, exhibit limited capability in capturing sufficiently fine-grained features, thereby constraining their effectiveness in identifying small and subtle defect targets, particularly in challenging transmission line inspection environments with interference from vegetation, man-made structures, and adjacent electrical installations.
Inspection of the heatmaps exported from ResNet further reveals that they lack clear edges and fine contours, which exacerbates the difficulty of accurately localizing subtle defects. Unlike conventional convolution, where the high performance of CNN-based edge feature extraction typically relies on large pretrained backbones—resulting in high memory consumption and energy cost, which contradicts the real-time requirements of UAV-based transmission line defect detection tasks—difference convolution introduces differential operations to enhance sensitivity to gradient variations and fine-grained structural details, thereby improving the extraction of local details and object boundariesWang et al. (2025).
To address the aforementioned problem, we propose an EEConv (edge enhancement convolution). Building on this idea, EEConv explicitly incorporates three difference convolution operators—central difference convolution (CDC), horizontal difference convolution (HDC), and vertical difference convolution (VDC)—to embed gradient priors into the feature extraction process. The overall structure of EEConv is illustrated in Fig. 2a.To ensure deployment efficiency, EEConv adopts a re-parameterization strategy that fuses its multi-branch convolutions into a single equivalent 3×3 kernel during inference. Once fused, the auxiliary branches are discarded, leaving only a lightweight standard convolution. This allows the module to preserve the expressive power gained from multi-branch training while maintaining the same computational cost as a vanilla convolution.During deployment, the structure of EEConv becomes as shown in Fig. 2b.
@@
Formally, let the three convolutional branches be denoted as $(W_{k},b_{k}),\;k=1,\dots,4$, corresponding to CDC, HDC, VDC and the conventional convolution, respectively. During training, their weights and biases are linearly combined:
$W_{\Sigma}$ $=\sum_{k=1}^{3}W_{k},$ (1)
$b_{\Sigma}$ $=\sum_{k=1}^{3}b_{k}.$ (2)
@@
To integrate this into our framework, we replace the second $3\times 3$ convolution in the residual branch of standard ResNet basic blocks with EEConv, thereby preserving the backbone’s modular design while substantially improving sensitivity to fine-grained structures. This design choice is motivated by the observation that the second $3\times 3$ convolution in each residual block primarily functions as a feature refiner, responsible for aggregating local structural details after the first convolution has extracted preliminary features. In this way, we obtain the Edge-Enhanced Block (EEBlock), whose structure is illustrated in Fig. 2c.
By substituting the refining convolution with EEConv, we embed directional gradient priors precisely at the stage where fine-grained information is consolidated, thus maximizing the impact of EEConv on enhancing edge and detail representation. Meanwhile, the first $3\times 3$ convolution and the shortcut connection remain unchanged, ensuring full compatibility with the original ResNet architecture and preserving its residual learning mechanism.
Integrating EEBlocks throughout the backbone yields the Edge-Enhanced ResNet (EE-ResNet), which significantly improves sensitivity to subtle structural details while maintaining computational efficiency. EE-ResNet provides a robust feature extraction foundation, offering stronger boundary-aware representations and improved resilience against background interference. This enhanced backbone design directly addresses the core challenge of detecting small-scale and fine-grained defects in transmission line imagery, thereby serving as a crucial cornerstone for the subsequent TinyDef-DETR framework. The overall architecture of the proposed EE-ResNet is illustrated in Fig. 3.
Figure 3: Overall framework of $E$-ResNet
@@ 3.4 Cross-Stage Dual-Domain Multi-Scale Attention Module
In transmission line defect detection, most targets are small-scale objects such as cracks, rust spots, or insulator damages. These defects often exhibit weak visual saliency and subtle structural variations, making them difficult to distinguish from complex backgrounds such as vegetation, sky, or towers. Conventional convolutional networks are limited by their local receptive fields, while large kernels may cause oversmoothing and incur high computational costs. Similarly, traditional attention mechanisms focusing solely on spatial or channel domains fail to provide sufficient enhancement for ambiguous and small targets. Therefore, a more effective mechanism is required to preserve fine-grained details, capture global semantics, and adaptively highlight defect-related features.
To address these challenges, we propose the Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM). Given input features $X\in\mathbb{R}^{B\times C\times H\times W}$, we first employ a $1\times 1$ convolution to adjust the channel dimension. The transformed features are then fed into parallel branches: a large-receptive-field branch, a local branch, and a dual-domain Frequency–Spatial–Channel Attention (FSCA) branch. The outputs are aggregated through residual summation and projected by a $1\times 1$ convolution to form the final response.To provide a clearer understanding of the proposed design, the overall architecture of CSDMAM is illustrated in Figure 5.
Figure 5: The architecture of Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM).
To effectively capture different types of information, we incorporate four complementary branches into the module.
Specifically, to enlarge the receptive field, we add a large-receptive-field branch employing depth-wise convolutions with kernel sizes $(k,k),(1,k),(k,1)$, where $k=31$. The square kernel models global contextual information, while the strip kernels capture elongated dependencies along horizontal and vertical directions, which are especially important for line-shaped structures in transmission infrastructure.
To preserve fine-grained local cues, we further introduce a local branch with a lightweight depth-wise $1\times 1$ convolution, ensuring that subtle defects such as cracks or corrosion spots are not smoothed out by large-kernel operations.
To strengthen semantic discrimination, we design a Frequency–Spatial–Channel Attention (FSCA) branch that jointly exploits frequency, spatial, and channel dependencies. Given input features $X\in\mathbb{R}^{B\times C\times H\times W}$, we first apply the two-dimensional Fourier transform to obtain the frequency representation
@@
where $\mathcal{F}$ denotes the discrete Fourier transform (DFT). Channel-wise weights are then computed via global average pooling followed by a $1\times 1$ convolution, producing $A_{c}\in\mathbb{R}^{B\times C\times 1}$. The enhanced frequency map is obtained as
$\tilde{X}_{f}=A_{c}\cdot X_{f},$
which is reconstructed into the spatial domain using the inverse DFT:
$X_{fs}=\mathcal{F}^{-1}(\tilde{X}_{f}).$
An additional spatial–channel reweighting map $A_{sc}$ is generated by global pooling and a $1\times 1$ convolution, yielding the FSCA-enhanced representation
$X_{fsca}=A_{sc}\cdot X_{fs}.$
The outputs of all branches are fused with the residual input to form the aggregated response:
$Y=X+DW_{1,k}(X_{n})+DW_{k,1}(X_{n})+DW_{k,k}(X_{n})+DW_{1,1}(X_{n})+X_{fsca},$
where $DW$ denotes depth-wise convolution with the specified kernel size. Finally, to improve efficiency and gradient propagation, we extend the design into a cross-stage structure. Following the Cross Stage Partial (CSP) strategy, the input is divided into two parts: one processed by the multi-branch attention operations and the other preserved as an identity mapping. The two parts are concatenated and fused through a $1\times 1$ convolution, giving the final output
$Y_{cs}=\text{Conv}([\text{DMAM}(X_{\text{branch}}),X_{\text{identity}}]).$
By combining large receptive fields, local detail preservation, and FSCA-based dual-domain enhancement within a cross-stage framework, CSDMAM achieves a favorable balance between global semantic understanding and precise defect localization, making it particularly effective for detecting small, visually ambiguous targets in UAV inspection imagery.